# Optical design and sensitivity optimization of Cryogenic sub-Hz cROss torsion bar detector with quantum NOn-demolition Speed meter (CHRONOS)

### v1

@@ VI.1 Noise model
The interferometer’s quantum noise is expressed in terms of the measured angular displacement $\theta$. Following the Buonanno-Chen formalism Buonanno and Chen (2003), we use the two-photon representation Caves and Schumaker (1981). In this method, we define the input and output quadrature vectors as
@@
where the input quadrature vector of the interferometer $\boldsymbol{a}$ is denoted. The amplitude quadrature (in-phase component) $a_{1}$ and the phase quadrature (out-of-phase component) $a_{2}$ are represented. The corresponding output quadrature vector $\boldsymbol{b}$ is also defined. With these definitions, the input-output relation can be written as
@@
where the transmittance of SRM $t_{s}$ and its phase delay $\Phi_{s}$ are obtained. The standard quantum limit (SQL) for angular displacement is Braginsky and Khalili (1992)
$\theta_{\mathrm{SQL}}=\sqrt{\frac{2\hbar}{I\Omega^{2}}}.$ (22)
@@
$\beta=\arctan\!\left(\frac{\Omega}{\gamma}\right),\qquad\beta_{\mathrm{sag}}=2\beta+\frac{\pi}{2}.$
@@
$\kappa_{\mathrm{sag}}(\Omega)=4\,\kappa(\Omega)\,|H_{\mathrm{PRC}}|^{2}\sin^{2}\beta_{\mathrm{sag}},$ (23)
where the frequency-dependent coupling $\kappa(\Omega)$ and cavity bandwidth $\gamma$ follow the definitions in Ref. Buonanno and Chen (2001). The PRC transfer function is
$H_{\mathrm{PRC}}=\frac{\sqrt{T_{p}}}{1-\sqrt{R_{p}}\,e^{2i\tau_{p}\Omega}r_{\mathrm{ifo}}},$
@@
where the one-way propagation time in the Sagnac loop $\tau_{\mathrm{sag}}$ and effective reflectance in ring cavities and Sagnac loop $r_{loop}$ are defined.
@@
$\theta=\frac{\theta_{\mathrm{SQL}}}{\sqrt{2\mathcal{K}_{\mathrm{sag}}(1-r_{s}^{2})}}\sqrt{\frac{\boldsymbol{\nu}^{\mathsf{T}}G\boldsymbol{\nu}}{\boldsymbol{\nu}^{\mathsf{T}}Q\boldsymbol{\nu}}},$ (24)
where the homodyne detection vector $\boldsymbol{\nu}^{\mathsf{T}}=(\cos\zeta,\,\sin\zeta)$ is defined. The $G$ and $Q$ matrices describe the amplitude-phase quadrature correlations induced by detuning and optomechanical coupling. Diagonalization yields an effective coupling constant $\kappa_{\mathrm{eff}}$ and an effective homodyne angle $\zeta_{\mathrm{eff}}$, from which the final quantum-limited sensitivity can be evaluated.
@@ VI.2 Optical Absorption and Mechanical Loss in Coatings
The optical coatings of the end test masses (ETMs) play a decisive role in setting both the quantum- and thermal-noise limits of the interferometer. Two major loss mechanisms—optical absorption and mechanical loss—must be simultaneously minimized to achieve the target sensitivity in the sub-hertz regime.
Optical absorption.
@@
Even ppm-level absorption leads to increased intracavity loss, reducing the effective reflectivity of the recycling cavities and lowering the circulating optical power Hello and Vinet (1990). Because both shot noise and radiation-pressure noise scale with optical power, quantum noise is highly sensitive to absorption: an increase in absorption simultaneously raises the shot-noise floor and reduces the radiation-pressure back-action, thereby rebalancing the overall quantum-noise spectrum.
@@
Figure 3 compares shot noise and radiation-pressure noise for different ETM coating reflectivities. The variation in shot noise is small, since it mainly depends on the circulating arm power, whereas radiation-pressure noise increases significantly at low frequencies due to incomplete back-action cancellation when the interferometer reflectivity $r_{\mathrm{ifo}}$ is reduced. This behavior is further illustrated in Fig. 4, which shows the degradation of quantum-nondemolition (QND) suppression under reduced $r_{\mathrm{ifo}}$.
@@
In CHRONOS, suppression of optical absorption is pursued through the development of advanced low-loss coatings. Plasma-enhanced chemical vapor deposition (PECVD) SiN/SiO2 and SiON multilayers have demonstrated mechanical-loss angles on the order of $10^{-5}$ at 100 Hz and room temperature Pan et al. (2018); Chao et al. (2023). However, their optical absorption remains higher than required for sub-hertz operation. Low-pressure chemical vapor deposition (LPCVD) coatings are currently being investigated as an alternative, offering the potential to reduce absorption while preserving low mechanical loss. A detailed study of LPCVD absorption properties is presented in V. Kumar et al. Kumar et al. (2025).
Mechanical loss.
@@
The mechanical-loss angle of the coating layers directly determines the amplitude of Brownian thermal fluctuations Levin (1998); Harry et al. (2002). At the 2.5 m CHRONOS scale, this coating Brownian noise defines the mid-frequency sensitivity floor, even under cryogenic operation. Experiments have shown that at $T\approx 10\penalty 10000\ \mathrm{K}$, typical ion-beam-sputtered (IBS) coatings exhibit loss peaks due to structural relaxation in the amorphous layers Yamamoto et al. (2006); Martin et al. (2010). By contrast, annealed SiON coatings demonstrate significantly lower loss at cryogenic temperatures, confirming their potential for next-generation cryogenic interferometers Pan et al. (2018); Chao et al. (2023).
@@
Figure 5 presents the total quantum-noise spectra, including shot noise, radiation-pressure noise, and other technical noise sources. In the 0.1-10 Hz band, where seismic noise dominates, the coating-related effects still determine the overall sensitivity. To achieve the CHRONOS sensitivity goal, the total effective coating loss angle must satisfy $\phi_{c}^{\mathrm{eff}}\leq 10^{-5}$ at cryogenic temperatures, while maintaining a reflectivity of $R\sim 99.9999\%$. This dual optimization—simultaneous reduction of optical absorption and mechanical loss— is essential to sustaining the high finesse ($\mathcal{F}\simeq 3.1\times 10^{4}$) and quantum-noise-limited performance required for the CHRONOS sub-hertz interferometer.
Figure 3: Comparison of shot noise and radiation-pressure noise for varying ETM reflectivity.
@@
Figure 5: Total quantum-noise spectra including shot noise, radiation-pressure noise, and other technical noise sources. Coating effects dominate the sensitivity in the 0.1-10 Hz band. The intensity noise is calculated elsewhere Tanabe et al. (2025).
@@ VI.3 Impact of power-recycling cavity detuning
The power-recycling cavity (PRC) plays a central role in enhancing the circulating optical power and shaping the quantum-noise spectrum of the interferometer. Importantly, the PRC reflectivity is not chosen to strictly satisfy resonance for the carrier field; instead, it is deliberately shifted from the exact resonance condition as a result of optimization against low-frequency losses and radiation-pressure noise. This approach follows the same philosophy as detuned signal-recycling interferometers Buonanno and Chen (2001), but here it is applied in a unique way to CHRONOS. This adjustment enables the interferometer to suppress sensitivity degradation at low frequencies while maintaining balanced performance across the detection band. Thus, the PRC is not only an element to enhance power, but also a strategic control parameter to optimize the distribution of quantum noise over the full observation band.
@@
Finally, Fig. 8 presents the total noise spectra including shot noise, radiation-pressure noise, and other technical noise sources such as seismic and suspension noise. It is evident that at the optimal point these contributions balance each other, providing the best sensitivity near 1 Hz. Away from this point, low-frequency radiation-pressure noise or high-frequency shot noise quickly dominates, leading to degraded sensitivity. This result highlights a unique aspect of CHRONOS: while conventional Michelson interferometers typically employ signal-recycling (SR) detuning to reconfigure their sensitivity curves Buonanno and Chen (2001), CHRONOS achieves sensitivity optimization primarily through PR detuning. Thus, PRC detuning represents a distinctive design feature of CHRONOS, enabling a new optimization strategy for suppressing quantum noise in the sub-hertz band.
@@
Figure 8: Total quantum-noise spectra including shot noise, radiation-pressure noise, and technical noise sources. The optimal point balances the three contributions, while detuning leads to degraded sensitivity. The intensity noise is calculated in elsewhere Tanabe et al. (2025)
@@ VI.4 Impact of signal-recycling cavity detuning
The signal-recycling mirror (SRM) serves as a key optical element that reshapes the interferometer’s frequency response by recycling the signal sidebands, thereby improving overall sensitivity. By tuning the reflectivity and detuning phase of the SRM, the signal-recycling cavity (SRC) can redistribute the detector’s response between low and high frequencies Buonanno and Chen (2001). However, in the CHRONOS configuration, detuning the SRC away from resonance does not improve the overall sensitivity.
Figure 9 shows the effect of SRC detuning on shot noise. At resonance ($\phi_{s}=0$), the signal sidebands are optimally transmitted to the output port, minimizing the shot-noise floor across a broad frequency range. When the SRC is detuned, however, the amplitude ($a_{1}$) and phase ($a_{2}$) quadratures in the two-photon formalism become mixed, causing a rotation of the optical field basis in quadrature space. This mixing corresponds to a 90 quadrature rotation around the optical phase space, which modifies the detected combination of amplitude and phase components. As a result, the shot noise changes uniformly across the frequency band, effectively shifting the entire sensitivity curve upward or downward depending on the detuning phase. Thus, SRC detuning does not enhance sensitivity in any particular band, but instead produces a uniform change in the overall shot-noise level.
Figure 10 illustrates the corresponding impact on radiation-pressure noise. In a Sagnac speed-meter interferometer, radiation-pressure noise exhibits a characteristic transition from $1/f^{2}$ to $1/f$ scaling at low frequencies. At resonance, the transition frequency is optimized so that radiation-pressure noise is minimized around 1 Hz. However, detuning the SRC alters the balance between the amplitude and phase quadratures, breaking the quantum non-demolition (QND) correlation that cancels back-action. Consequently, radiation-pressure noise increases at low frequencies, degrading the sub-hertz sensitivity that is critical for CHRONOS.
@@
It is worth noting that in longer-baseline interferometers, such as the 300 m-scale system studied by Inoue et al. Inoue et al. (2025), a small optimal detuning of approximately $2^{\circ}$ has been identified. In that regime, the suppression of coating-induced time-delay effects becomes sufficient, and a small SRC detuning can slightly enhance the sensitivity. However, for the 2.5 m CHRONOS prototype, maintaining the exact resonance condition is most effective for preserving QND back-action cancellation and achieving quantum-noise-limited operation.
@@
Figure 11: Total quantum-noise spectrum including shot noise, radiation-pressure noise, and technical noise. At resonance, the three contributions balance near 1–10 Hz, while detuning leads to significant degradation. The intensity noise is calculated in elsewhere Tanabe et al. (2025)
@@ VI.5 Impact of Homodyne Detection Angle
The homodyne detection angle $\zeta$ determines the quadrature of the output field that is measured in the balanced homodyne detector (BHD). By adjusting $\zeta$, the interferometer effectively changes the relative weighting between the amplitude ($a_{1}$) and phase ($a_{2}$) quadratures in the readout. This tuning allows the measured signal to emphasize either the shot-noise or radiation-pressure-noise contribution, depending on the chosen detection quadrature Buonanno and Chen (2001); Danilishin et al. (2019).
@@
Figure 13 illustrates the corresponding radiation-pressure noise. At $\zeta=90^{\circ}$, the interferometer operates as an ideal speed meter, coherently cancelling radiation-pressure noise in the mid-frequency band. At lower frequencies, however, the QND condition is partially broken due to coating-induced time-delay effects in the end test masses (ETMs), and the residual $1/f^{2}$ component becomes visible. The transition frequency between the QND (back-action evasion) regime and the $1/f^{2}$ region depends sensitively on the homodyne angle $\zeta$. By fine-tuning $\zeta$, the interferometer can extend the frequency range over which the QND cancellation remains effective, thereby broadening the low-noise region of CHRONOS.
The combined quantum-noise spectra are shown in Fig. 14. The optimal homodyne angle is found to be $\zeta_{\mathrm{opt}}\simeq 46^{\circ}$, where the trade-off between radiation-pressure suppression and shot-noise minimization yields the best overall sensitivity around 1 Hz. This configuration also ensures robust operation against optical gain fluctuations. Angles significantly smaller than $40^{\circ}$ or larger than $60^{\circ}$ result in degraded performance, demonstrating that precise control of the homodyne phase is essential for maintaining quantum-noise-limited operation in CHRONOS.
@@
Figure 14: Total quantum-noise spectra for various homodyne angles. The optimal configuration at $\zeta_{\mathrm{opt}}\simeq 46^{\circ}$ achieves balanced suppression of radiation-pressure and shot noise, yielding the best sensitivity near 1 Hz. The intensity noise is calculated in elsewhere Tanabe et al. (2025).

### v2

@@ VI.1 Noise model
The interferometer’s quantum noise is expressed in terms of the measured angular displacement $\theta$. Following the Buonanno–Chen formalism Buonanno2003 , we use the two-photon representation Caves1981 . In this method, we define the input and output quadrature vectors as
@@
Here, the input quadrature vector of the interferometer is denoted by $\boldsymbol{a}$. The amplitude quadrature (in-phase component) $a_{1}$ and the phase quadrature (out-of-phase component) $a_{2}$ are represented. The corresponding output quadrature vector $\boldsymbol{b}$ is also defined. With these definitions, the input–output relation can be written as
@@
where the transmittance of the SRM is $t_{s}=\sqrt{1-R_{s}}$ and its phase delay is $\Phi_{s}$. We define $G=CC^{\mathsf{T}}$, and $Q=\Re(\boldsymbol{D})\,\Re(\boldsymbol{D})^{\mathsf{T}}+\Im(\boldsymbol{D})\,\Im(\boldsymbol{D})^{\mathsf{T}}$, where the matrix $C=[c_{ij}]$ and the column vector $\boldsymbol{D}^{\mathsf{T}}=(D_{1},D_{2})$ are defined following Eqs. (2.22)–(2.24) of Ref. BuonannoChen2001 . The standard quantum limit (SQL) for angular displacement is BraginskyKhalili1992
$\theta_{\mathrm{SQL}}(\Omega)=\sqrt{\frac{2\hbar}{I\Omega^{2}}}.$ (27)
@@
$\beta(\Omega)=\arctan\!\left(\frac{\Omega}{\gamma}\right),\qquad\beta_{\mathrm{sag}}(\Omega)=2\beta(\Omega)+\frac{\pi}{2}.$
@@
$\mathcal{K}_{\mathrm{sag}}(\Omega)=4\,\mathcal{K}(\Omega)\,|H_{\mathrm{PRC}}(\Omega)|^{2}\sin^{2}\beta_{\mathrm{sag}}(\Omega),$ (28)
where the frequency-dependent coupling $\mathcal{K}(\Omega)$ and cavity bandwidth $\gamma$ follow the definitions in Ref. BuonannoChen2001 . The PRC transfer function is
$H_{\mathrm{PRC}}(\Omega)=\frac{\sqrt{1-R_{p}}}{1-\sqrt{R_{p}}\,e^{2i\tau_{p}\Omega}r_{\mathrm{ifo}}},$
@@
where the one-way propagation time in the Sagnac loop $\tau_{\mathrm{sag}}$ and the effective reflectance in the ring cavities and Sagnac loop $r_{\mathrm{loop}}\sim r_{i}$ are defined.
@@
$\theta(\Omega)=\frac{\theta_{\mathrm{SQL}}(\Omega)}{\sqrt{2\mathcal{K}_{\mathrm{sag}}(\Omega)(1-r_{s}^{2})}}\sqrt{\frac{\boldsymbol{\nu}^{\mathsf{T}}G\boldsymbol{\nu}}{\boldsymbol{\nu}^{\mathsf{T}}Q\boldsymbol{\nu}}},$ (29)
where the homodyne detection vector $\boldsymbol{\nu}^{\mathsf{T}}=(\cos\zeta,\,\sin\zeta)$ is defined. Diagonalization yields an effective coupling constant $\mathcal{K}_{\mathrm{eff}}$ and an effective homodyne angle $\zeta_{\mathrm{eff}}$. The effective homodyne angle is obtained as $\zeta_{\mathrm{eff}}=\zeta-\psi(f)+\theta(f)$, where $\psi(f)$ and $\theta(f)$ denote the principal-axis rotations of $G$ and $Q$, respectively.
@@ VI.2 Optical Absorption and Mechanical Loss in Coatings
The optical coatings of the end test masses play a decisive role in setting both the quantum- and thermal-noise limits of the interferometer. Two major loss mechanisms—optical absorption and mechanical loss—must be simultaneously minimized to achieve the target sensitivity in the sub-hertz regime.
Optical absorption
@@
Even ppm-level absorption leads to increased intracavity loss, reducing the effective reflectivity of the recycling cavities and lowering the circulating optical power Hello1990 . Because both shot noise and radiation-pressure noise scale with optical power, quantum noise is highly sensitive to absorption: an increase in absorption simultaneously raises the shot-noise floor and reduces the radiation-pressure back-action, thereby rebalancing the overall quantum-noise spectrum.
@@
Figures 3 and 4 compare shot noise and radiation-pressure noise for different ETM coating reflectivities. The variation in shot noise is small, since it mainly depends on the circulating arm power, whereas radiation-pressure noise increases significantly at low frequencies due to incomplete back-action cancellation when the interferometer reflectivity $r_{\mathrm{ifo}}$ is reduced. This behavior is further illustrated in Fig. 4, which shows the degradation of QND suppression under reduced $r_{\mathrm{ifo}}$.
@@
In CHRONOS, suppression of optical absorption is pursued through the development of advanced low-loss coatings. Plasma-enhanced chemical vapor deposition (PECVD) SiN/SiO2 and SiON multilayers have demonstrated mechanical-loss angles on the order of $10^{-5}$ at 100 Hz and room temperature Pan2018_SiN_Chao ; Chao2023_SiON . However, their optical absorption remains higher than required for sub-hertz operation. Low-pressure chemical vapor deposition (LPCVD) coatings are currently being investigated as an alternative, offering the potential to reduce absorption while preserving low mechanical loss. A detailed study of LPCVD absorption properties is presented in V. Kumar et al. Vivek:inprep .
Mechanical loss
@@
The mechanical-loss angle of the coating layers directly determines the amplitude of Brownian thermal fluctuations Levin1998 ; Harry2002 . At the 2.5 m CHRONOS scale, this coating Brownian noise defines the mid-frequency sensitivity floor, even under cryogenic operation. Experiments have shown that at $T\approx 10\penalty 10000\ \mathrm{K}$, typical ion-beam-sputtered (IBS) coatings exhibit loss peaks due to structural relaxation in the amorphous layers Yamamoto2006 ; Martin2010 . By contrast, annealed SiON coatings demonstrate significantly lower loss at cryogenic temperatures, confirming their potential for next-generation cryogenic interferometers Pan2018_SiN_Chao ; Chao2023_SiON .
@@
Figure 5 presents the total quantum-noise spectra, including shot noise, radiation-pressure noise, and other technical noise sources. In the 0.1–10 Hz band, where seismic noise dominates, the coating-related effects still determine the overall sensitivity. To achieve the CHRONOS sensitivity goal, the total effective coating loss angle must satisfy $\phi_{c}^{\mathrm{eff}}\leq 10^{-5}$ at cryogenic temperatures, while maintaining a reflectivity of $R\sim 99.9999\%$. This dual optimization—simultaneous reduction of optical absorption and mechanical loss— is essential to sustaining the high finesse ($\mathcal{F}\simeq 3.1\times 10^{4}$) and quantum-noise-limited performance required for the CHRONOS sub-hertz interferometer.
Figure 3: Impact of shot noise for varying ETM reflectivity.
@@
Figure 5: Total quantum-noise spectra including shot noise, radiation-pressure noise, and other technical noise sources. Coating effects dominate the sensitivity in the 0.1–10 Hz band. The intensity noise is calculated elsewhere Tanabe:inprep .
@@ VI.3 Impact of power-recycling cavity detuning
The PRC plays a central role in enhancing the circulating optical power and shaping the quantum-noise spectrum of the interferometer. Importantly, the PRC reflectivity is not chosen to strictly satisfy resonance for the carrier field; instead, it is deliberately shifted from the exact resonance condition as a result of optimization against low-frequency losses and radiation-pressure noise. This approach follows the same philosophy as detuned signal-recycling interferometers BuonannoChen2001 , but here it is applied in a unique way to CHRONOS. This adjustment enables the interferometer to suppress sensitivity degradation at low frequencies while maintaining balanced performance across the detection band. Thus, the PRC is not only an element to enhance power, but also a strategic control parameter to optimize the distribution of quantum noise over the full observation band.
@@
Finally, Fig. 8 presents the total noise spectra including shot noise, radiation-pressure noise, and other technical noise sources such as seismic and suspension noise. It is evident that at the optimal point these contributions balance each other, providing the best sensitivity near 1 Hz. Away from this point, low-frequency radiation-pressure noise or high-frequency shot noise quickly dominates, leading to degraded sensitivity. This result highlights a unique aspect of CHRONOS: while conventional Michelson interferometers typically employ SR detuning to reconfigure their sensitivity curves BuonannoChen2001 , CHRONOS achieves sensitivity optimization primarily through PR detuning. Thus, PRC detuning represents a distinctive design feature of CHRONOS, enabling a new optimization strategy for suppressing quantum noise in the sub-hertz band.
@@
Figure 8: Total quantum-noise spectra including shot noise, radiation-pressure noise, and technical noise sources. The optimal point balances the three contributions, while detuning leads to degraded sensitivity. The intensity noise is calculated elsewhere Tanabe:inprep .
@@ VI.4 Impact of signal-recycling cavity detuning
The SRM serves as a key optical element that reshapes the interferometer’s frequency response by recycling the signal sidebands, thereby improving overall sensitivity. By tuning the reflectivity and detuning phase of the SRM, the SRC can redistribute the detector’s response between low and high frequencies BuonannoChen2001 . However, in the CHRONOS configuration, detuning the SRC away from resonance does not improve the overall sensitivity.
Figure 9 shows the effect of SRC detuning on shot noise. At resonance ($\phi_{s}=0$), the signal sidebands are optimally transmitted to the output port, minimizing the shot-noise floor across a broad frequency range. When the SRC is detuned, however, the amplitude ($a_{1}$) and phase ($a_{2}$) quadratures in the two-photon formalism become mixed, causing a rotation of the optical field basis in quadrature space. This mixing corresponds to a $90^{\circ}$ quadrature rotation in optical phase space, which modifies the detected combination of amplitude and phase components. As a result, the shot noise changes uniformly across the frequency band, effectively shifting the entire sensitivity curve upward or downward depending on the detuning phase. Thus, SRC detuning does not enhance sensitivity in any particular band, but instead produces a uniform change in the overall shot-noise level.
Figure 10 illustrates the corresponding impact on radiation-pressure noise. In a Sagnac speed-meter interferometer, radiation-pressure noise exhibits a characteristic transition from $1/f^{2}$ to $1/f$ scaling at low frequencies. At resonance, the transition frequency is optimized so that radiation-pressure noise is minimized around 1 Hz. However, detuning the SRC alters the balance between the amplitude and phase quadratures, breaking the QND correlation that cancels back-action. Consequently, radiation-pressure noise increases at low frequencies, degrading the sub-hertz sensitivity that is critical for CHRONOS.
@@
It is worth noting that in longer-baseline interferometers, such as the 300 m-scale system studied by Inoue et al. inoue2025chronos , a small optimal detuning of approximately $2^{\circ}$ has been identified. In that regime, the suppression of coating-induced time-delay effects becomes sufficient, and a small SRC detuning can slightly enhance the sensitivity. However, for the 2.5 m CHRONOS prototype, maintaining the exact resonance condition is most effective for preserving QND back-action cancellation and achieving quantum-noise-limited operation.
@@
Figure 11: Total quantum-noise spectrum including shot noise, radiation-pressure noise, and technical noise. At resonance, the three contributions balance near 1–10 Hz, while detuning leads to significant degradation. The intensity noise is calculated elsewhere Tanabe:inprep .
@@ VI.5 Impact of Homodyne Detection Angle
The homodyne detection angle $\zeta$ determines the quadrature of the output field that is measured in the balanced homodyne detector. By adjusting $\zeta$, the interferometer effectively changes the relative weighting between the amplitude ($a_{1}$) and phase ($a_{2}$) quadratures in the readout. This tuning allows the measured signal to emphasize either the shot-noise or radiation-pressure-noise contribution, depending on the chosen detection quadrature BuonannoChen2001 ; Danilishin2019 .
@@
Figure 13 illustrates the corresponding radiation-pressure noise. At $\zeta=90^{\circ}$, the interferometer operates as an ideal speed meter, coherently cancelling radiation-pressure noise in the mid-frequency band. At lower frequencies, however, the QND condition is partially broken due to coating-induced time-delay effects in the ETMs, and the residual $1/f^{2}$ component becomes visible. The transition frequency between the QND regime and the $1/f^{2}$ region depends sensitively on the homodyne angle $\zeta$. By fine-tuning $\zeta$, the interferometer can extend the frequency range over which the QND cancellation remains effective, thereby broadening the low-noise region of CHRONOS.
The combined quantum-noise spectra are shown in Fig. 14. The optimal homodyne angle is found to be $\zeta\simeq 46^{\circ}$, where the trade-off between radiation-pressure suppression and shot-noise minimization yields the best overall sensitivity around 1 Hz. This configuration also ensures robust operation against optical gain fluctuations. Angles significantly smaller than $40^{\circ}$ or larger than $60^{\circ}$ result in degraded performance, demonstrating that precise control of the homodyne phase is essential for maintaining quantum-noise-limited operation in CHRONOS.
@@
Figure 14: Total quantum-noise spectra for various homodyne angles. The optimal configuration at $\zeta\simeq 46^{\circ}$ achieves balanced suppression of radiation-pressure and shot noise, yielding the best sensitivity near 1 Hz. The intensity noise is calculated elsewhere Tanabe:inprep .